A rabbit matures at 6 months, and has a gestation period of 30 days (call it one month) with a 4 week time lapse after birth to become fertile again (call it one month). We will assume that we start with one breeding pair of rabbits and that each litter contains 10 rabbits, half of which are female. Such a population of rabbits increases as shown in the table below. Month: 0 1 2 3 4 5 6 7 8 9 10 11 12 Population: 2 12 12 22 22 32 32 92 92 202 202 362 362 14. How would you make a decision about which type of function to use to model a population such as this population of rabbits?
Answer #1

exponential regression

y = a b^x

ln y = ln a + x ln b

running regression of ln y on x

Month Population ln y
0 2 0.693147
1 12 2.484907
2 12 2.484907
3 22 3.091042
4 22 3.091042
5 32 3.465736
6 32 3.465736
7 92 4.521789
8 92 4.521789
9 202 5.308268
10 202 5.308268
11 362 5.891644
12 362 5.891644
SUMMARY OUTPUT
Regression Statistics
Multiple R 0.970939707
R Square 0.942723915
Adjusted R Square 0.937516998
Standard Error 0.386101875
Observations 13
ANOVA
df SS MS F Significance F
Regression 1 26.99029948 26.99029948 181.0522313 3.55662E-08
Residual 11 1.639821239 0.149074658
Total 12 28.63012072
Coefficients Standard Error t Stat P-value Lower 95%
Intercept 1.55249915 0.202372371 7.671497571 9.70982E-06 1.107080565
Month 0.385095246 0.028619775 13.45556507 3.55662E-08 0.322103546

ln y = 1.5525 + 0.3851 x

y = e^(1.5525 + 0.3851 x)

= 4.723263 * 1.46975 ^x